### v2

@@ 5.1. A Counterexample of Monotonicity
where $M(t)=K\,t+L\sqrt{1-t^{2}}$. Let $\|Q_{A}\|=\|Q_{B}\|=1$ and $\rho_{A}=\rho_{B}$ such that $\cos(\rho_{A})=\cos(\rho_{B})=0.6$. Since the angle between $Q_{A}$ and $Q_{B}$ is $2\rho_{A}$, the law of cosines implies that $\|Q\|=(2\|Q_{A}\|^{2}-2\|Q_{A}\|^{2}\cos(\pi-2\rho_{A}))^{1/2}=1.2$. Then, by choosing $z_{1}=(x_{1},y_{1})=(0.49,0.1)$ and $z_{2}=(x_{2},y_{2})=(0.1,0.94)$, we derive that $F(z_{1})\approx(0.998,-0.111)$ and $F(z_{2})\approx(0.675,0.324)$, hence $F(z_{1})-F(z_{2})\approx(0.323,-0.436)$ and $z_{1}-z_{2}=(0.39,-0.84)$. We obtain that $(F(z_{1})-F(z_{2}))^{\top}(z_{1}-z_{2})>0$. Let $\bm{\theta_{i}}=(\arccos(x_{i}),\arccos(y_{i}))$ for $i=1,2$. Note that by $\mathop{}\!\mathrm{d}(\cos\theta)/\mathop{}\!\mathrm{d}(\theta)=-\sin(\theta)$, the pseudo-gradient $\tilde{F}$ with respect to $\theta_{A}$ and $\theta_{B}$ is $(F_{1}(x,y)\cdot(-\sqrt{1-x^{2}}),F_{2}(x,y)\cdot(-\sqrt{1-y^{2}}))$, we still have $(\tilde{F}(\bm{\theta}_{1})-\tilde{F}(\bm{\theta}_{2}))^{\top}(\bm{\theta}_{1}-\bm{\theta}_{2})>0$. Hence, this is a counterexample of monotonicity which also disproves the $\lambda$-cocoercivity of the game for any $\lambda>0$.
@@ 6. Discretization Comes to the Rescue
Algorithm 1 Grid-Based Approx. $\varepsilon$-PSNE Search (GBA-PSNE)
@@
2:for $i\leftarrow 0$ to N1N-1 do
3:  $\mathbf{x}_{i}\leftarrow$ rotate $Q_{A}/\|Q_{A}\|$ toward $Q$ for angle $\min\{ih,\rho_{A}\}$, then append $\mathbf{x}_{i}$ to $G_{A}$
4:  $\mathbf{y}_{i}\leftarrow$ rotate $Q_{B}/\|Q_{B}\|$ toward $Q$ for angle $\min\{ih,\rho_{B}\}$, then append $\mathbf{y}_{i}$ to $G_{B}$
5:end for
6:for each $\mathbf{y}\in G_{B}$ do
7:  $\hat{\mathbf{x}}\leftarrow\max_{\mathbf{x}\in G_{A}}R_{A}(\mathbf{x},\mathbf{y})$,   $\hat{\mathbf{y}}\leftarrow\max_{\mathbf{y}\in G_{B}}R_{B}(\hat{\mathbf{x}},\mathbf{y})$
[…]
10:  if $\text{gain}_{A}\leq\hat{\varepsilon}$ and $\text{gain}_{B}\leq\hat{\varepsilon}$ then
[…]
12:  end if
13:end for
Algorithm 2 Ternary Best Response (TBR)
@@
(sketch) Given the grid $G_{A}\times G_{B}$ from the algorithm, with a slight abuse of notation, let $(a_{i},b_{j})\in G_{A}\times G_{B}$ be a strategy profile that is an $\hat{\varepsilon}-$PSNE on the grid such that $R_{A}(a_{i},b_{j}):=f_{\theta_{A}}(a_{i},b_{j})\geq\max_{a_{\ell}\in G}f_{\theta_{A}}(a_{\ell},b_{j})-\hat{\varepsilon}\mbox{ and }R_{B}(a_{i},b_{j}):=g_{\theta_{B}}(a_{i},b_{j})\allowbreak\geq\max_{b_{\ell}\in G}g_{\theta_{B}}(a_{i},b_{\ell})-\hat{\varepsilon}$. Since $h\leq\frac{\varepsilon}{2\mathcal{L}}$ and $\hat{\varepsilon}=\varepsilon-\mathcal{L}h$, we can derive that $(a_{i},b_{j})$ is an $\varepsilon$-PSNE in the game. The time complexity of the algorithm can be shown to be $O(nk+k/\varepsilon^{2})$ (see Appendix F). In addition, since $R_{A}$ and $R_{B}$ are one-dimensional quasi-concave functions for fixed $r_{A}=r_{B}=1$, they are unimodal so that Line 7 of Algorithm GBA-PSNE can be replaced by $\hat{\mathbf{x}}\leftarrow\texttt{TBR}(G_{A},[0,N-1],R_{A}(\cdot,\mathbf{y}))$ and $)\hat{\mathbf{y}}\leftarrow\texttt{TBR}(G_{B},[0,N-1],R_{B}(\hat{\mathbf{x}}$$\cdot$))”. The original grid-search time $O(N^{2})=O(1/\epsilon^{2})$ can then be shrunk to $O(N\log N)=O((1/\epsilon)\log(1/\epsilon))$.

### v3

@@ 5. Bad and Good News for the Gradient-Based Approach
Remark
Throughout this section we use the pseudo-gradient of payoffs, $F$ is the concatenated pseudo-gradients of players’ payoffs, not their costs. Consequently, the associated VI is written with $\leq$”. One can equivalently work with F-F and the standard $\geq$” VI convention.
@@ 5.1. A Counterexample of Monotonicity
where $M(t)=K\,t+L\sqrt{1-t^{2}}$. Let $\|Q_{A}\|=\|Q_{B}\|=1$ and $\rho_{A}=\rho_{B}$ such that $\cos(\rho_{A})=\cos(\rho_{B})=0.6$. Since the angle between $Q_{A}$ and $Q_{B}$ is $2\rho_{A}$, the law of cosines implies that $\|Q\|=(2\|Q_{A}\|^{2}-2\|Q_{A}\|^{2}\cos(\pi-2\rho_{A}))^{1/2}=1.2$. Then, by choosing $\mathbf{v}_{1}=(x_{1},y_{1})=(0.40,0)$ and $\mathbf{v}_{2}=(x_{2},y_{2})=(0,0.43)$, we derive that $F(\mathbf{v}_{1})\approx(0.5049,0.4049)$ and $F(\mathbf{v}_{2})\approx(0.4058,0.5096)$, hence $F(\mathbf{v}_{1})-F(\mathbf{v}_{2})\approx(0.0991,-0.1047)$ and $\mathbf{v}_{1}-\mathbf{v}_{2}=(0.40,-0.43)$. We obtain that $(F(z_{1})-F(z_{2}))^{\top}(z_{1}-z_{2})\approx 0.0847>0$. Let $\bm{\theta_{i}}=(\arccos(x_{i}),\arccos(y_{i}))$ for $i=1,2$. Note that by $\mathop{}\!\mathrm{d}(\cos\theta)/\mathop{}\!\mathrm{d}(\theta)=-\sin(\theta)$, the pseudo-gradient $\tilde{F}$ with respect to $\theta_{A}$ and $\theta_{B}$ is $(F_{1}(x,y)\cdot(-\sqrt{1-x^{2}}),F_{2}(x,y)\cdot(-\sqrt{1-y^{2}}))$, we still have $(\tilde{F}(\bm{\theta}_{1})-\tilde{F}(\bm{\theta}_{2}))^{\top}(\bm{\theta}_{1}-\bm{\theta}_{2})<0$. Hence, this is a counterexample of monotonicity which also disproves the $\lambda$-cocoercivity of the game for any $\lambda>0$.
@@ 6. Discretization Comes to the Rescue
Algorithm Grid-Based Approx. $\varepsilon$-PSNE Search (GBA-PSNE)
[…]
0: a grid profile that is an $\varepsilon$-PSNE of the continuous game
1: Compute the Lipschitz constant $\mathcal{L}$ for $f_{\theta_{A}}$ and $g_{\theta_{B}}$.
2: Set the grid spacing $h$ so that $h\leq\varepsilon/(4\mathcal{L})$, and let $\hat{\varepsilon}\leftarrow 3\mathcal{L}h$.
3: Construct the uniform grids with spacing $h$
$G_{A}=\{a_{0},a_{1},\ldots,a_{N-1}\}\subseteq[0,\rho_{A}],$
$G_{B}=\{b_{0},b_{1},\ldots,b_{N-1}\}\subseteq[0,\rho_{B}].$
4:for $j=0$ to N1N-1 do
5:  Compute $M_{A}(j):=\max_{0\leq i\leq N-1}f_{\theta_{A}}(a_{i},b_{j})$;
6:  Compute the interval $I_{A}(j)$ of all indices $i$ satisfying $f_{\theta_{A}}(a_{i},b_{j})\geq M_{A}(j)-\hat{\varepsilon}$.
7:end for
8:for $i=0$ to N1N-1 do
9:  Compute $M_{B}(i):=\max_{0\leq j\leq N-1}g_{\theta_{B}}(a_{i},b_{j})$;
10:  Compute the interval $I_{B}(i)$ of all indices $j$ satisfying $g_{\theta_{B}}(a_{i},b_{j})\geq M_{B}(i)-\hat{\varepsilon}$.
11:end for
12: Run Algorithm Sweep-Intervals on the family of intervals $\{I_{A}(j)\}_{j=0}^{N-1}$ and $\{I_{B}(i)\}_{i=0}^{N-1}$.
13: Let $(i,j)$ be the returned pair and return $(a_{i},b_{j})$
@@
0: a pair $(i,j)$ such that $i\in I_{A}(j)$ and $j\in I_{B}(i)$, if any.
1: Initialize empty lists $\mathrm{Add}[0],\ldots,\mathrm{Add}[N-1]$ and $\mathrm{Rem}[0],\ldots,\mathrm{Rem}[N-1]$.
2:for $i=0$ to N1N-1 do
3:  Append $i$ to $\text{Add}[I^{|\leftarrow}_{B}(i)]$.
4:  if $I^{\rightarrow|}_{B}(i)+1\leq N-1$ then
5:   Append $i$ to $\text{Rem}[I^{\rightarrow|}_{B}(i)+1]$.
6:  end if
7:end for
8: Initialize an empty ordered set $\mathcal{A}$.
9:for $j=0$ to N1N-1 do
10:  for $i\in\text{Rem}[j]$ do
11:   delete $i$ from $\mathcal{A}$.
12:  end for
13:  for $i\in\text{Add}[j]$ do
14:   insert $i$ into $\mathcal{A}$.
15:  end for
16:  /* $\mathcal{A}=\{i\mid j\in I_{B}(i)\}$: the set of active indices at step $j$ */
17:  Let $t\leftarrow\min\{i\in\mathcal{A}\mid i\geq I^{|\leftarrow}_{A}(j)\}$.
18:  if $t\neq\text{null}$ and $t\leq I^{\rightarrow|}_{A}(j)$ then
19:   return $(t,j)$
20:  end if
21:end for
22:return $\emptyset$ /* “no pair found” */
Algorithm Ternary Best Response (TBR)
@@
(sketch) Given the grid $G_{A}\times G_{B}$ from the algorithm, with a slight abuse of notation, let $(a_{i},b_{j})\in G_{A}\times G_{B}$ be a strategy profile that is an $\hat{\varepsilon}-$PSNE on the grid such that $R_{A}(a_{i},b_{j}):=f_{\theta_{A}}(a_{i},b_{j})\geq\max_{a_{\ell}\in G}f_{\theta_{A}}(a_{\ell},b_{j})-\hat{\varepsilon}\mbox{ and }R_{B}(a_{i},b_{j}):=g_{\theta_{B}}(a_{i},b_{j})\allowbreak\geq\max_{b_{\ell}\in G}g_{\theta_{B}}(a_{i},b_{\ell})-\hat{\varepsilon}$. Since $h\leq\frac{\varepsilon}{2\mathcal{L}}$ and $\hat{\varepsilon}=\varepsilon-\mathcal{L}h$, we can derive that $(a_{i},b_{j})$ is an $\varepsilon$-PSNE in the game. The time complexity of the algorithm can be shown to be $O(nk+k/\varepsilon^{2})$. In addition, since $R_{A}$ and $R_{B}$ are one-dimensional quasi-concave functions for fixed $r_{A}=r_{B}=1$, they are unimodal so that Lines 5,6,9,10 of Algorithm GBA-PSNE can be computed using Algorithm TBR for the ternary search and binary searches for the endpoints of intervals of best-responses. The original grid-search time $O(N^{2})=O(1/\epsilon^{2})$ can then be shrunk to $O(N\log N)=O((1/\epsilon)\log(1/\epsilon))$.